OpenStudy (anonymous):

true or false two cities are 15.7 cm apart on a map that has a scale of 1cm : 14 km. these cities are 211.4 km apart.

OpenStudy (anonymous):

Dimensional analysis:\[15.7 cm \times \frac{14 km}{1 cm}\]yields an answer in kilometers.

OpenStudy (anonymous):

TRUE OR FALSE?

OpenStudy (anonymous):

Perform the calculation and see how far apart they actually are. Then compare that to the number they gave.

OpenStudy (anonymous):

OK THANKS

OpenStudy (anonymous):

Here's another way to look at it. \[\frac {1cm}{14km} = \frac{15.7cm}{x km}\]